40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 15 12:25:38 2011, kiwamu, Update, Green Locking, Y arm ALS : beat-note free run fluctuation noise_budget.png
    Reply  Wed Nov 16 17:29:46 2011, kiwamu, Update, Green Locking, Some updates on the Y end green PDH Yarm_fluctuation.png
       Reply  Thu Nov 17 17:21:22 2011, kiwamu, Update, Green Locking, Y end green PDH servo : it's okay servoTF.pngservofilter.pngerr_suppression.png
          Reply  Fri Nov 18 11:28:39 2011, rana, Update, Green Locking, Y end green PDH servo : it's okay 
             Reply  Fri Nov 18 12:11:24 2011, Zach, Update, Green Locking, Y end green PDH servo : it's okay 
    Reply  Thu Nov 17 18:20:26 2011, kiwamu, Update, Green Locking, Y arm ALS : 1st trial of noise budget Yarm_ALS_2011Nov16.png
       Reply  Mon Nov 21 16:08:04 2011, kiwamu, Update, Green Locking, 2nd trial of Y arm ALS noise budget : broad band noise gone Yarm_ALS_2011Nov19.png
          Reply  Tue Nov 22 15:18:18 2011, kiwamu, Update, Green Locking, broad band noise depends on the gain of Y green PDH. and comaprator broken 
Message ID: 5930     Entry time: Thu Nov 17 18:20:26 2011     In reply to: 5894     Reply to this: 5970
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y arm ALS : 1st trial of noise budget 

The noise budget on the Y arm ALS has begun.

Right now the fluctuation of the green beat-note seems mostly covered by unknown noise which is relatively white.

(Though I feel I made a wrong calibration ... I have to check it again)

 

Yarm_ALS_2011Nov16.png

(Measurement condition)

 + The Y arm is locked to the PSL laser by acting o ETMY.
 + The end green laser is locked to the Y arm.
 + The fine resolution MFD (Mixer-base Frequency Discriminator) is used to observe the beat-note fluctuation
   (We have two MFDs : fine resolution and coarse resolution.)
ELOG V3.1.3-